# Language Models, Graph Searching, and Supervision Adulteration: When More Supervision is Less and How to Make More More

### v1

@@ B.11 Future Token Prediction
Wu et al. (2024b) studied the mechanism for LMs to learn future information from a next-token prediction objective. They hypothesized that it could be due to two mechanisms; a deliberate pre-cashing mechanism which computes features earlier than they are needed and an unintentional breadcrumb mechanism which considers that a LM learns features for predicting the next token and that these just happen to also be good features for predicting future tokens also. They construct a synthetic dataset and show that pre-cashing is done and necessary for some planning tasks. However, they also show that pre-cashing is less noticeable in a GPT2 model used for natural language (but also consider this might be less true as LMs scale up). Notably this mechanism will not help for the path-star task as future predictions can ignore both re-cashing and breadcrumb features due to the CHC. That is, any features used for planning will just be ignored. This also means that there will be no learning signal to reinforce learning such features.
We design future distributions and associated losses to enhance this ability for the path-star task. Not only do these create an explicit belief-state that allows for planning, they also avoid adulteration by targeting tokens that require multiple edges or path-reconstruction to predict. RITF is also designed to be efficient during train as it only requires a single parallelizable loss. This is in contrast to other losses on $N$ tokens into the future, which scale linearly with $N$ (Goodman et al., 2020; Qi et al., 2020).

### v3

@@ C.11 Future Token Prediction
Wu et al. (2024b) studied the mechanism for LMs to learn future information from a next-token prediction objective. They hypothesized that it could be due to two mechanisms; a deliberate pre-cashing mechanism which computes features earlier than they are needed and an unintentional breadcrumb mechanism which considers that a LM learns features for predicting the next token and that these just happen to also be good features for predicting future tokens also. They construct a synthetic dataset and show that pre-cashing is done and necessary for some planning tasks. However, they also show that pre-cashing is less noticeable in a GPT2 model used for natural language (but also consider this might be less true as LMs scale up). Notably this mechanism will not help for the path-star task as future predictions can ignore both pre-cashing and breadcrumb features due to the CHC. That is, any features used for planning will just be ignored. This also means that there will be no learning signal to reinforce learning such features.
We design future distributions and associated losses to enhance this ability for the path-star task. Not only do these create an explicit belief-state that allows for planning, they also avoid adulteration by targeting tokens that require multiple edges or path-reconstruction to predict. RITF is also designed to be efficient during train as it only requires a single parallelizable loss. This is in contrast to other losses on $N$ tokens into the future, which scale linearly with $N$ (Goodman et al., 2020; Qi et al., 2020; Gloeckle et al., 2024).